OpenStudy (anonymous):

Calculate the derivative of the function. g(x) = (x^2 − 3x − 1)^−4 ?

OpenStudy (saifoo.khan):

-4(x^2-3x-1)^-5 (2x-3)

OpenStudy (saifoo.khan):

\[\frac{-4(2x-3)}{x^2-3x-1}\]

OpenStudy (anonymous):

you forgot that the denominator is to the power of 5

OpenStudy (anonymous):

\[\frac{12-8 x}{\left(x^2-3 x-1\right)^5} \]